# ACoT-VLA: Action Chain-of-Thought for Vision-Language-Action Models

### v1

@@ Appendix C More Experimental Results
In this section, we provide additional quantitative experiments to substantiate the effectiveness of our proposed approach and to empirically uncover several insightful phenomena. Specifically, the experimental analyses comprise three parts: (1) ablation study conducted on the LIBERO-Plus benchmark in Table 9, (2) an investigation of how the parameter sizes of the Action Head and Explicit Action Reasoner (EAR), as well as the number of denoising steps, influence policy performance in Table 10, and (3) a comparative study examining the relationship among inference latency, model size, and performance in Table 11. Note that we adopt $\pi_{0.5}$ [22] as the baseline method, denoted as “Baseline”.
@@
Table 11: Ablation experiment on model efficiency and performance.

### v2

@@ Appendix C More Experimental Results
In this section, we provide additional quantitative experiments to substantiate the effectiveness of our proposed approach and to empirically uncover several insightful phenomena. Specifically, the experimental analyses comprise four parts: (1) ablation study conducted on the LIBERO-Plus benchmark in Table 9, (2) an investigation of how the parameter sizes of the Action Head and Explicit Action Reasoner (EAR), as well as the number of denoising steps, influence policy performance in Table 10, (3) evaluation of proposed approach’s sim2real capability based on Genie-Sim 3.0 [53] in Table 11, and (4) a comparative study examining the relationship among inference latency, model size, and performance in Table 12. Note that we adopt $\pi_{0.5}$ [22] as the baseline method, denoted as “Baseline”.
@@
Tasks Simulation Real-World
$\pi_{0.5}$ Ours $\pi_{0.5}$ Ours
Select Color 86.0 98.8 85.0 94.0
Recognize Size 93.0 96.0 94.0 94.0
Grasp Targets 71.7 68.0 70.8 75.0
Organize Objects 52.0 74.0 60.0 68.4
Avg. 75.7 84.2 77.5 82.9
Table 11: Experimental results on Genie Sim 3.0 Simulation and Real-world Transfer. The best results are highlighted in bold.
Sim To Real Evaluation. To further assess the sim2real transferability of proposed method, we conduct evaluations on the Genie-Sim 3.0 benchmark [53]. It comprises $4$ challenging tasks: picking objects based on specified colors, sizes, or categories, as well as a complex tabletop organization task. Technically, the model is trained on the officially released simulation-based datasets. Then, it is deployed and evaluated in both simulation and real-world environments.
As illustrated in Table 11, our approach consistently outperforms the baseline across both domains. Specifically, it achieves $84.2\%$ in simulation and $82.9\%$ in real-world settings, representing absolute improvements of $8.5\%$ and $5.4\%$, respectively. Notably, our method exhibits minimal performance degradation during the sim-to-real transition, which we attribute to the fundamental nature of ACoT paradigm. Specifically, while visual domain gap persists between synthetic and real-world observations, the underlying kinematically grounded action guidance remains consistent. By shifting the locus of reasoning from the perceptual space to the action space, our model extracts task-relevant motion priors that are invariant to low-level visual perturbations, effectively enhancing policy’s sim2real capability.
@@
Table 12: Ablation experiment on model efficiency and performance. Note that the evaluation protocol in LIBERO-Plus is Supervised Fine-Tuning.